Johnny is designing a rectangular poster to contain 32 inches squared of printing with a 4​-in margin at the top and bottom and a 2​-in margin at each side. What overall dimensions will minimize the amount of paper​ used?
1
Expert's answer
2020-04-30T17:48:16-0400

Consider the printed area of horizontal length "x" and vertical length "y" as shown in the figure below:





Area of printed region as shown above in the figure is,



"xy=32"


"y=\frac{32}{x}"

Here, the objective is to minimize the total area of the poster"A=(x+4)(y+8)"


Substitute "\frac{32}{x}" for "y" into "A=(x+4)(y+8)" to express the total area in terms of single variable "x" as,



"A(x)=(x+4)(\frac{32}{x}+8)"

Differentiate with respect to "x" and equate to zero in order to find the critical value as,


"A'(x)=0"


"(x+4)\frac{d}{dx}(\frac{32}{x}+8)+(\frac{32}{x}+8)\frac{d}{dx}(x+4)=0"


"(x+4)(-\frac{32}{x^2})+(\frac{32}{x}+8)(1)=0"


"-\frac{32}{x}-\frac{128}{x^2}+\frac{32}{x}+8=0"


"-\frac{128}{x^2}+8=0"


"x^2=16"


"x=4"


Next plug "x=4" into "y=\frac{32}{x}" to obtain "y=\frac{32}{4}=8"


Here, the first derivative change sign from negative to positive at "x=4" , therefore, there exists a minima at "x=4" .


Hence, the overall dimensions of poster is:
Horizontal length of poster"=x+4=4+4=8" in.Vertical length of poster"=y+8=8+8=16" in.